At auction, 5-month T-bills were sold at a discount of 7.430%. What was the simple annual yield r? HINT [See Example 5.] (Round your answer to three decimal places.)
Answer #1

Solution : Computing the sample annual gield of the bond, we have - 5 month T-bills were sold at a discount of 7.430% . → Let= (1000/92517) 126 1 = (1.20367) - 1 = 0.20357 X 100 = 20.357 %. Hence, The of the bond is Sample annual field 20:367 %